1.determine the molecular formula of compounds from
a)a certain gaseous hydrocarbon contains 80%carbon and 20%hydrogen by mass 224cm³of the compound at stp 0.3g
b)a certain gaseous compound contains 42.86% and 57.14% oxygen 100cm³of the compound at stp weighs 0.125g
1
Expert's answer
2020-04-10T13:35:39-0400

a) C:H = 80/12 : 20/1 = 1:3

Formula is (CH3)n , molar mass M = 15n g/mole,

at stp m (g) / M (g/mole) = V (cm3)/22400 (cm3/mole)

M = 0.3/224*22400 = 30 g/mole

n = 30/15 = 2, (CH3)2

compound is C2H6

b) C:O = 42.86/12 : 57.14/16 = 1 : 1

formula is (CO)n

molar mass is M = 28n

at stp M = 0.125*22400/100 = 28 g/mole

compound is CO